Write the following problem as a common fraction or mixed number. Write your answer in the lowest terms.
15.35
Thanks

Found 2 solutions by benni1013, josmiceli:
Answer by benni1013(206)

15.35==>15 35/100==>15 7/20

Answer by josmiceli(19441)

15.35
.35 is 35/100
as a mixed number:
15.35 = 15 35/100
15 35/100 = 15 7/20
as a fraction:
15 = (20/20) * 15
15 = 300/20
15 7/20 = 300/20 + 7/20 = 307/20